OpenStudy (anonymous):

What is the measure of the hypotenuse? 17 289 169 13

OpenStudy (anonymous):



OpenStudy (raden):

use the pythagorean theorem : x^2 = 5^2 + 12^2 simplify, then solve for x

OpenStudy (anonymous):

144 + 25 = x^2 169 = xˆ2 169 / 2 ??

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

x^2 = 169 to get x, take the squareroot of 169 so, what is sqrt(169) ?

OpenStudy (anonymous):

srry 13

OpenStudy (raden):

that;s right